# On the existence of unbiased resilient estimators in discrete quantum systems

### v2

@@ III.1 Classical
We consider a family of N-point support probability functions $P_{\theta}(x)$ where $P_{\theta}(x_{i})\geq 0$ for $x_{i}\in\{x_{1},..,x_{N}\}$. The family of probability distributions is parametrized by the parameter $\theta$ that we would like to infer.
@@
Since we are considering discrete probability distributions the mean value can be written as finite sum $\langle\tilde{\Theta}\rangle_{\theta_{0}}=\sum_{i}P_{\theta_{0}}(x_{i})\tilde{\Theta}(x_{i})$. Defining $\tilde{\Theta}(x_{i})=\tilde{\theta_{i}}$, Eq. (23) can be written as a system of equations $A\vec{x}=\vec{b}$ where
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \\ \end{pmatrix},\quad\quad\vec{x}=\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$ (24)
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}|_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution. If $n>N-1$, then the higher derivatives are no longer independent, and they can be written as a combination of the first $N-1$ linearly independent derivatives. In this case, the system has no solution or the solution is trivial, that is to say, the $n^{th}$order BhB is equal to $N-1$ order BhB for all $n>N-1$. It is then concluded that given a $N$ points probability distribution is advisable to compute BhB bound up to order $N-1$.
In the appendix (B), the conditions for the existence of the estimator are given. It is also shown that the bound diverges if and only if the estimator does not exist.
@@ III.2 Quantum
We extend the previous results to determine when computing QBhB gives more information about the estimation problem. In quantum estimation theory the measurement is a POVM $\{\Pi_{i}\}$, where $i$ are the possible outcomes, each one with probability $P_{\theta}(i)=\operatorname{Tr}(\Pi_{i}\rho_{\theta})$. The classical processing of the measurement outcomes is accounted by the estimator $\tilde{\Theta}(x)$ that associates an estimate $\tilde{\Theta}(i)$ to the outcome $i$. Every estimate $\tilde{\Theta}(i)$ has a probability given by $P_{\theta}(i)$. Both tasks, measurement, and classical processing, can be treated simultaneously as a hermitian operator $\hat{\Theta}=\sum_{i}\tilde{\Theta}(i)\Pi_{i}$ [22]. The conditions of Eq. (23) are now written in terms of the operator $\hat{\Theta}$ as
@@
where we are using the Einstein summation convention and $i,j=1,2,...,N$. Now we have a system of equations with complex coefficients for the complex variables $\hat{\Theta}_{ij}$. Since the estimator operator must be hermitian, not all the variables $\hat{\Theta}_{ij}$ are independent. Instead of $N^{2}$ independent variables, there are now $\frac{N(N+1)}{2}$ independent variables. In that way, we can vectorize the system of equations as follows
$\begin{pmatrix}(\rho_{\theta_{0}})_{1}&(\rho_{\theta_{0}})_{2}&\dots&(\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d\rho_{\theta_{0}})_{1}&(d\rho_{\theta_{0}})_{2}&\dots&(d\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d^{2}\rho_{\theta})_{1}&(d^{2}\rho_{\theta})_{2}&\dots&(d^{2}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \vdots&\vdots&\ddots&\vdots\\ (d^{n}\rho_{\theta})_{1}&(d^{n}\rho_{\theta})_{2}&\dots&(d^{n}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \end{pmatrix}\begin{pmatrix}\hat{\Theta}_{1}\\ \hat{\Theta}_{2}\\ \vdots\\ \hat{\Theta}_{\frac{N(N+1))}{2}}\end{pmatrix}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix}.$ (47)
@@
So we have a system of equations equivalent to the one for a classical probability distribution with $\frac{N(N+1)}{2}$ support points. As before if this system of equations cannot be solved the QBhB variance will go to infinity. On the other hand, if the first $\frac{N(N+1)}{2}-1$ derivatives are linearly independent, for all $n>\frac{N(N+1)}{2}-1$ one has that $n^{th}$order QBhB is equal to $\frac{N(N+1)}{2}-1$ order QBhB. The difference between the quantum and classical cases is that in the former we have the freedom to choose a measurement. This extra freedom implies that the highest relevant QBhB bound –which gives a different bound than lower orders– is lifted in the quantum case because $N-1<\frac{N(N+1)}{2}-1$.
@@
Let us elaborate more on this because at first glance the result may seem to be at odds with the classical case. After projective measurements, one deals with $N$ point probability distributions. Therefore, the results derived in the classical setting must also apply, so it might seem that there will be a contradiction to what we stated here. To solve this apparent contradiction, we explain why $N$ order QBh bound need not be equal to the $N-1$ one. The key point is that to compute the $N$ and the $N+1$ order QBhB, different measurements have to be performed, giving different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the measurements is what gives the quantum systems more non-trivial information, or in other words, the maximal order for which QBhB gives different information is higher than for classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in section (III.1).
III.3 Example of the quantum case
In this section, we show a case in which we see the lifting in the constraints produced by the quantum nature of the problem, meaning that we see that the quantum bounds are not trivial –giving the same value as bounds from lower orders– for $n>N-1$, being $n$ the order of the BhB employed and $N$ the number of outcomes.
Consider that our state $\rho(\theta)$ is obtained through the following unitary operation $U=e^{-i\theta^{2}H}$ where $H=\sigma_{x}$ is the $x$-Pauli matrix, and we would like to estimate $\theta$. The number of outcomes here is $N=2$. Our initial state is given by
$\rho(0)=\frac{1}{2}\mathbb{I}+\frac{(2\lambda-1)}{2}\sigma_{z},$ (48)
where $\lambda\in(0,1)$ and
IV Conclusions
@@
Since long ago, it has been clear in estimation theory that while the CRB is extremely useful and illustrative is not always meaningful. For that reason, different bounds were constructed and introduced in the classical regime. On the other hand, until very recently, there were no works extending these bounds to the quantum regime [30].
In this article, we first motivated the use of Bhattacharyya bounds using a simple example. Essentially, when one does not know the prior value of the parameter exactly, but knows that is within a small range, the BhB are more desirable. We then gave necessary and sufficient conditions for an estimator that satisfies the BhB to exist. The existence conditions are related to the support of the probability distribution which determines the freedom we have to construct an estimator. When we computed the $n$-th order BhB bound we found a limit for the variance of an estimator that satisfies $n+1$ unbiasedness conditions. When the support of the probability distribution is small, we might not have enough freedom to construct an estimator satisfying the desired constraints. Namely, the estimator exists if and only if the maximization presented in Eq. (8) can be calculated (non-divergent). Moreover, following a similar reasoning, we could anticipate when computing higher orders in the BhB gives no extra information. Specifically, we saw that for $N$ outcomes, it is “sufficient” to calculate the $n^{th}$order BhB bound where $n\leq N-1$. However, in the quantum case, where we have more freedom in the measurements, the order is lifted to $n\leq N(N+1)/2-1$ implying that in the quantum regime, the estimators are more resilient. Finally, it would be of theoretical and practical interest to extend the results presented here to continuous probability distributions as well as to biased estimators. For the latter case, we think it will be conceptually difficult to compare the different bounds such as the BhB and CRB.
@@ Acknowledgements.
We acknowledge fruitful discussions with Manuel Gessner and Maximilian Reichert. The authors acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, as well as from the EU FET-Open project EPIQUS (Grant No. 899368), also from Project Grant No. PID2021-125823NA-I00 595 and Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future,” this project has also received support from the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call –Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan– NextGenerationEU within the framework of the Digital Spain 2026 Agenda, we acknowledge funding from Basque Government through Grant No. IT1470-22 and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government.
Appendix B Existence of the estimator
@@
B.1 Bhattacharyya bounds for pure states
In this section, we show that contrary to what was recently claimed in the literature [30], Bhattacharyya bounds for pure states are computable and can be different from the Cramer-Rao bound in general. For this aim, we first explore why the conclusion exposed in [30], namely that for pure states Bhattacharyya bounds are the same as Cramer-Rao bound is faulted. After that, we present a simple counterexample of the conclusion. The proof in [30] is based on the fact that for pure states, $L_{i}$ (which represent the analogous of SLD but for higher derivatives) are bounded. From this, the authors concluded that the Bhattacharyya matrix is then filled with unbounded terms and when taking the optimization problem, one only looks at the first entry of the matrix which corresponds to the quantum Fisher Information. As exposed in [Yating], even when the operators $L_{i}$ are unbounded the product of $\rho L_{i}L_{j}$ may not be so, and this is what appears in the BhBattacarya matrix. The authors there argue that the usual Cramer-Rao bound (and for extension all Bhattacharyya bounds) is only constructed for bounded operators. For that reason, they “rederive” the Cramer-Rao bound for a newly defined operator $Q=L\sqrt{\rho}$. In the same fashion, one could do the same with Bhattacharyya bounds and write the Bhattacharyya matrix as a function of $Q_{i}$ operators. [RR: This we have to explicitly do here. ]This is shown in the next example. Consider a state $|\psi\rangle=\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$ where $x$ is the parameter one wants to estimate. From this state one can construct $Q_{i}$ operators that fulfill
$\frac{1}{2}(\sqrt{\rho}Q^{\dagger}_{j}+Q_{j}\sqrt{\rho})=\frac{\partial^{j}\rho}{\partial x^{j}},$ (97)
$\sqrt{\rho}Q_{j}-Q^{\dagger}_{j}\sqrt{\rho}=0,$ (98)
where the second equation comes from the hermiticity of $L_{j}$ [Yating]. After simple algebra one gets that (up to second order in Bhattacharyya bounds) in the basis $\{|\psi\rangle,|\psi^{\perp}\rangle\}$ the operators read
$Q_{1}=\begin{pmatrix}\langle\psi|\frac{\partial\rho}{\partial x}|\psi\rangle&0\\ 2\langle\psi^{\perp}|\frac{\partial\rho}{\partial x}|\psi\rangle&0\end{pmatrix},$ (99)

### v3

@@ III.1 Classical
We consider a family of $N$-point support probability distributions $P_{\theta}(x)$ where $P_{\theta}(x_{i})\geq 0$ for $x_{i}\in\{x_{1},..,x_{N}\}$. The family of probability distributions is parametrized by the parameter $\theta$ that we would like to infer.
@@
Since we are considering discrete probability distributions the mean value can be written as finite sum $\langle\tilde{\Theta}\rangle_{\theta_{0}}=\sum_{i}P_{\theta_{0}}(x_{i})\tilde{\Theta}(x_{i})$. Defining $\tilde{\Theta}(x_{i})=\tilde{\theta_{i}}$, Eq. (LABEL:condiciones) can be written as a system of equations $A\vec{x}=\vec{b}$ where
$A=\begin{pmatrix}P_{\theta_{0}}(x_{1})&P_{\theta_{0}}(x_{2})&\dots&1-\sum_{i}^{N-1}P_{\theta_{0}}(x_{i})\\ \partial P_{\theta_{0}}(x_{1})&\partial P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial P_{\theta_{0}}(x_{i})\\ \partial^{2}P_{\theta_{0}}(x_{1})&\partial^{2}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{2}P_{\theta_{0}}(x_{i})\\ \vdots&\vdots&\ddots&\vdots\\ \partial^{n}P_{\theta_{0}}(x_{1})&\partial^{n}P_{\theta_{0}}(x_{2})&\dots&-\sum_{i}^{N-1}\partial^{n}P_{\theta_{0}}(x_{i})\\ \\ \end{pmatrix}$
and
$\vec{x}=$ $\begin{pmatrix}\tilde{\theta}_{1}\\ \tilde{\theta}_{2}\\ \vdots\\ \tilde{\theta}_{N}\end{pmatrix},\quad\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$
where we defined $\partial^{l}P_{\theta_{0}}(x)\equiv\frac{\partial^{l}P_{\theta}(x)}{\partial\theta^{l}}\big{|}_{\theta=\theta_{0}}$. The system has $n+1$ equations and $N$ variables. We consider the case where the first $N-1$ derivatives are linearly independent. This implies that if $n\leq N-1$, there is always a solution.
If $n>N-1$, then the higher derivatives are no longer independent and can be written as a combination of the first $N-1$ linearly independent derivatives. In this case, the system has no solution, or the solution is trivial. That is to say, the $n^{\text{th}}$-order BhB is equal to the $(N-1)$-order BhB for all $n>N-1$.
It is then concluded that for a probability distribution with $N$ points, it is advisable to compute the BhB bound up to order $N-1$. In the appendix (A), the conditions for the existence of the estimator are given. It is also shown that the bound diverges if and only if the estimator does not exist.
@@ III.2 Quantum
We extend the previous results to determine when computing QBhB gives more information about the estimation problem. In quantum estimation theory the measurement is a POVM $\{\Pi_{i}\}$, where $i$ are the possible outcomes, each one with probability $P_{\theta}(i)=\operatorname{Tr}(\Pi_{i}\rho_{\theta})$. The classical processing of the measurement outcomes is accounted by the estimator $\tilde{\Theta}(x)$ that associates an estimate $\tilde{\Theta}(i)$ to the outcome $i$. Every estimate $\tilde{\Theta}(i)$ has a probability given by $P_{\theta}(i)$. Both tasks, measurement, and classical processing, can be treated simultaneously as a hermitian operator $\hat{\Theta}=\sum_{i}\tilde{\Theta}(i)\Pi_{i}$ [23]. The conditions of Eq. (LABEL:condiciones) are now written in terms of the operator $\hat{\Theta}$ as
@@
where we are using the Einstein summation convention and $i,j=1,2,...,N$ where $N$ is the dimension of the Hilbert space of the physical system considered. Now we have a system of equations with complex coefficients for the complex variables $\hat{\Theta}_{ij}$. Since the estimator operator must be hermitian, not all the variables $\hat{\Theta}_{ij}$ are independent. Instead of $N^{2}$ independent variables, there are now $\frac{N(N+1)}{2}$ independent variables. The system of equations can be vectorized and rewritten as $A\vec{x}=\vec{b}$ where,
$A=\begin{pmatrix}(\rho_{\theta_{0}})_{1}&(\rho_{\theta_{0}})_{2}&\dots&(\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d\rho_{\theta_{0}})_{1}&(d\rho_{\theta_{0}})_{2}&\dots&(d\rho_{\theta_{0}})_{\frac{N(N+1)}{2}}\\ (d^{2}\rho_{\theta})_{1}&(d^{2}\rho_{\theta})_{2}&\dots&(d^{2}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \vdots&\vdots&\ddots&\vdots\\ (d^{n}\rho_{\theta})_{1}&(d^{n}\rho_{\theta})_{2}&\dots&(d^{n}\rho_{\theta})_{\frac{N(N+1)}{2}}\\ \end{pmatrix}$ (21)
and
$\vec{x}=$ $\begin{pmatrix}\hat{\Theta}_{1}\\ \hat{\Theta}_{2}\\ \vdots\\ \hat{\Theta}_{\frac{N(N+1))}{2}}\end{pmatrix},\quad\quad\quad\vec{b}=\begin{pmatrix}\theta_{0}\\ 1\\ 0\\ \vdots\\ 0\end{pmatrix},$
@@
So we have a system of equations equivalent to the one for a classical probability distribution with $\frac{N(N+1)}{2}$ support points. As before if this system of equations cannot be solved the QBhB variance will go to infinity. On the other hand, if the first $\frac{N(N+1)}{2}-1$ derivatives are linearly independent, for all $n>\frac{N(N+1)}{2}-1$ one has that $n^{th}$order QBhB is equal to $\frac{N(N+1)}{2}-1$ order QBhB. The difference between the quantum and classical cases is that in the former we have the freedom to choose a measurement. This extra freedom implies that the highest relevant QBhB bound –which gives a different bound than lower orders– is lifted in the quantum case because $N-1<\frac{N(N+1)}{2}-1$. In appendix (B) an example is studied in detail.
@@
Let us elaborate further on this because, at first glance, the result may seem to be at odds with the classical case. After projective measurements, one deals with $N$-point probability distributions. Therefore, the results derived in the classical setting must also apply, and it might seem that there is a contradiction to what we stated here.
To resolve this apparent contradiction, we explain why the $N$-order quantum Bhattacharyya bound (QBhB) need not be equal to the $(N-1)$-order QBhB. The key point is that, to compute the $N$- and $(N+1)$-order QBhB, different measurements have to be performed, resulting in different probability distributions $P_{N}$ and $P_{N+1}$. This extra freedom in the choice of measurements allows quantum systems to access more non-trivial information. In other words, the maximal order for which the QBhB provides distinct information is higher than in classical scenarios. Thus, since $P_{N}$ and $P_{N+1}$ are different probability distributions, there is no discrepancy with what we presented in Section (III.1).
IV Mach-Zehnder interferometer
Historically, greater attention has been given to the Barankin bounds [27, 28]. In contrast, to the best of our knowledge, Bhattacharyya bounds have not been computed for any physical problem before. In this work, we compute the Bhattacharyya bounds and identify the estimators that achieve these bounds for the paradigmatic Mach-Zehnder interferometer. We demonstrate that, when the prior distribution is not sharply peaked and we move away from the point where unbiased estimators are constructed, the Bhattacharyya bounds provide an improvement in the achievable precision.
$\varphi_{a}$$\varphi_{b}$$a$$b$$a^{\prime}$$b^{\prime}$|$\ket{r}$|$\ket{r}$$n_{a}$$n_{b}$
Figure 2: Mach-Zehnder interferometer setup. A 50:50 beam splitter is illuminated by two spatially separate input modes, represented by the annihilation operators $\hat{a}$ and $\hat{b}$, respectively. Both input modes are initialized in identical Fock states $|r\rangle$, with $r$ denoting the fixed photon number in each mode. Upon interaction with the first beam splitter, the input modes are coherently mixed, producing two output modes. These output modes then propagate through paths experiencing distinct phase shifts, denoted by $\varphi_{a}$ and $\varphi_{b}$. The phase-modulated modes are subsequently recombined at a second 50:50 beam splitter. Finally, photon number measurements are performed on the output modes of the second beam splitter, enabling interference-based observations of the system.
IV.0.1 Mach-Zehnder
In the Mach-Zehnder configuration, depicted in Fig. (2), two modes of the electromagnetic field interfere on a balanced beamsplitter, the output beams acquire a relative phase $\theta=\varphi_{a}-\varphi_{b}$, and interfere again on the second beamsplitter. Finally, the photon numbers $n_{a}$ and $n_{b}$ are measured at the output ports.
We consider the input state |$\ket{r,r}$, where both input modes are in the Fock state with $r$ photons. It has been shown that this state achieves the Heisenberg limit [38] while estimating the relative phase $\theta$. The evolution induced by the interferometer is given by the unitary operator $U=e^{\frac{\theta}{2}(a^{\dagger}b-b^{\dagger}a)}$ [6]. The probability distribution $P_{\theta}(2q)$ is then
$P_{\theta}(2q)=J^{2}_{q}(r\theta),$ (23)
where $J_{q}(x)$ are the Bessel functions. We compute the CRB, the $2^{\text{nd}}$-order BhB, and the estimators achieving each bound at $\theta_{0}=10^{-3}$. In Fig. (3), we compare the mean squared error (MSE) of these two estimators within an interval of $\theta$ that contains $\theta_{0}$.
One can conclude that when the prior knowledge is not sharp, the BhB offers a more robust bound; i.e., $\text{MSE}_{\text{Bh}}$ remains flat over an interval of $\theta$, whereas $\text{MSE}_{\text{CR}}$ increases rapidly when $\theta\neq\theta_{0}$. In addition, the protocol that saturates the BhB achieves a smaller MSE. Therefore, it may be advisable to use this bound as a figure of merit instead of the CRB in certain specific scenarios.
(a) MSE
(b) Bias
Figure 3: Results for the input state |$\ket{r,r}$ for $r=5000$ and $\theta_{0}=0.001$.
In Fig. (4), we analytically compute the second-order Bhattacharyya bound and the Cramér-Rao bound as a function of the increasing number of photons. We observe that the Cramér-Rao bound achieves the Heisenberg limit, as expected. Interestingly, the Bhattacharyya bound also reaches the Heisenberg limit, indicating that a tighter bound can still achieve Heisenberg scaling. This result suggests an alternative approach to constructing estimators that can achieve optimal scaling even with limited prior knowledge of the true parameter value. In particular, this demonstrates that it is possible to design estimation protocols that are more robust to uncertainties in prior knowledge while still reaching the Heisenberg limit in precision.
Figure 4: A logarithmic plot of the Cramér-Rao bound and the second-order Bhattacharyya bound is shown. Both bounds scale as $\sim\frac{1}{n^{2}}$, which means that both bounds attain the Heisenberg limit.
V Conclusions
@@
Since long ago, it has been clear in estimation theory that while the Cramér-Rao bound (CRB) is extremely useful and illustrative, it is not always meaningful. For that reason, different bounds were constructed and introduced in the classical regime. On the other hand, until very recently, there were no works extending these bounds to the quantum regime [30].
In this article, we first motivated the use of Bhattacharyya bounds by studying its properties in several examples. Essentially, when one does not know the prior value of the parameter exactly, but knows that it is within a small range, the BhB are more desirable. We then gave necessary and sufficient conditions for an estimator that satisfies the BhB to exist. The existence conditions are related to the support of the probability distribution, which determines the freedom we have to construct an estimator.
When we computed the $n$-th order Bhattacharyya bound, we found a limit for the variance of an estimator that satisfies $n+1$ unbiasedness conditions. When the support of the probability distribution is small, we might not have enough freedom to construct an estimator satisfying the desired constraints. Namely, the estimator exists if and only if the maximization presented in Eq. (6) can be calculated (non-divergent).
Moreover, following a similar reasoning, we could anticipate when computing higher orders in the BhB gives no extra information. Specifically, we saw that for $N$ outcomes, it is “sufficient” to calculate the $n$-th order BhB bound where $n\leq N-1$. However, in the quantum case, where we have more freedom in the measurements, the order is lifted to $n\leq\frac{N(N+1)}{2}-1$, implying that in the quantum regime, the estimators are more resilient. Finally, we compute the Bhattacharyya bound for a Mach-Zehnder interferometer and numerically show that this bound scales inversely proportional to the square of the number of photons, which is the Heisenberg scaling. Since Bhattacharyya bounds are tighter, this result supports the claim that it is possible to achieve Heisenberg scaling in the Mach-Zehnder interferometer using a more robust estimator.
We note that the existence of quantum Barankin bounds for finite dimensional systems has also been studied recently in Ref. [39].
@@ Acknowledgements.
We acknowledge fruitful discussions with Manuel Gessner and Maximilian Reichert. The authors acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, as well as from the EU FET-Open project EPIQUS (Grant No. 899368), also from Project Grant No. PID2021-125823NA-I00 595 and Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future,” this project has also received support from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026, and by the EU through the Recovery, Transformation and Resilience Plan– NextGenerationEU within the framework of the Digital Spain 2026 Agenda, we acknowledge funding from Basque Government through Grant No. IT1470-22 and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government.
Appendix A Existence of the estimator
@@
Appendix B Example of the quantum case
In this section, we show a case in which we see the lifting in the constraints produced by the quantum nature of the problem, meaning that we see that the quantum bounds are not trivial –giving the same value as bounds from lower orders– for $n>N-1$, being $n$ the order of the BhB employed and $N$ the number of measurement outcomes. As in this example we deal with qubits, $N=2$.
Consider that our state $\rho(\theta)$ is obtained through the following unitary operation $U=e^{-i\theta^{2}H}$ where $H=\sigma_{x}$ is the $x$-Pauli matrix, and we would like to estimate $\theta$. The initial state is given by
$\rho(0)=\frac{1}{2}\mathbb{I}+\frac{(2\lambda-1)}{2}\sigma_{z},$ (39)
where $\lambda\in(0,1)$ and
Now we consider the classical probability distributions that arise after performing a measurement. An optimal measurement to achieve the QCRB consists of a projection onto the eigenstates of $L_{1}$. On the other hand, a measurement that saturates the $n$-th order BhB is given by a projection onto the eigenstates of
$\alpha(Q^{-1}\bm{\lambda})\bm{L},$
where $\bm{L}=(L_{1},L_{2},\ldots,L_{n})$, $Q$ is the BhB matrix, $\bm{\lambda}=(1,0,..,0)$ and $\alpha$ is a normalization constant [30]. The resulting probability distributions obtained by measuring the state $\rho_{\theta}$ with the optimal measurement—which saturates the CRB in the point $\phi$ are:
$P_{\theta}^{\text{CRB}}(\alpha_{1})$ $=\frac{(2\lambda-1)\sin f-1}{\left(\sin\left(2\phi^{2}\right)-1\right)\left(\left(\tan\left(2\phi^{2}\right)+\sec\left(2\phi^{2}\right)\right)^{2}+1\right)}$
$P_{\theta}^{\text{CRB}}(\alpha_{2})$ $=1-P_{\theta}^{\text{CRB}}(\alpha_{1})$
where $f=2(\theta-\phi)(\theta+\phi)$ and $\alpha_{i}$ are the eigenvalues of $L_{1}$.
The probability distribution that saturates the BhB is,
$P_{\theta}^{Bh}(\beta_{1})=$ $\frac{4\phi^{2}\big{(}(2\lambda-1)\left(\sin\left(2\theta^{2}-4\phi^{2}\right)-g\sin f\right)+\sin\left(2\phi^{2}\right)\left(4(1-2\lambda)\phi^{2}\sin f+g\right)\big{)}+(1-2\lambda)g\cos f}{32\phi^{4}+2g\left(4\phi^{2}\sin\left(2\phi^{2}\right)-\cos\left(2\phi^{2}\right)\right)+2}$
$+\frac{\cos\left(2\phi^{2}\right)\left((2\lambda-1)\cos f-g\right)+16\phi^{4}+1}{32\phi^{4}+2g\left(4\phi^{2}\sin\left(2\phi^{2}\right)-\cos\left(2\phi^{2}\right)\right)+2}$
$P_{\theta}^{Bh}(\beta_{2})=$ $1-P_{\theta}^{Bh}(\beta_{1})$
where $g=\sqrt{16\phi^{4}+1}$ and $\beta_{i}$ are eigenvalues of $\alpha(Q^{-1}\bm{\lambda})\bm{L}$. Now we compute the BhB matrix for $P_{\theta}^{\text{CRB}}(x)$ and for $P_{\theta}^{Bh}(x)$ obtaining,
$\bm{C}_{\text{CRB}}=\left(\begin{array}[]{cc}16\theta^{2}(1-2\lambda)^{2}&16\theta(1-2\lambda)^{2}\\ 16\theta(1-2\lambda)^{2}&16(1-2\lambda)^{2}\\ \end{array}\right)$ (42)